how to find the wavelength of the radiation when the ionization energy of hydrogen is iven


 

Grade:

2 Answers

Godfrey Classic Prince
12 years ago

Dear abi ram,

It is simple to find the Wavelength of the Radiation if the Ionisation Energy of Hydrogen is given ..

I.E = 13.6 * Z2/n2

Now here Z=1 ,Since it is a Hydrogen atom !

n is the energy level in which it is present !!

But according to you the Value of Ionisation Energy is given so , you have the I.E Value !

Now, Ionisation Energy = hc/λ where h is the planks constant & c is the velocity of the light !!

hc = 19.8 * 10-26

Now you can easily get the Value of λ(Wavelength) ...

 

Hope thishelped you immensely..!

If you loved the answer then please do not forget to approve it by clicking on "Yes" below.!!Smile

chandrakiran anem
12 years ago

basically       E=hv    v can be written in terms of wavelenth

 

       wavelength=frequency(v)/sppeed of light